OpenStudy (anonymous):

this is a pythagorean theorem im trying to find c? A= 3 B= 9 what is C?

jhonyy9 (jhonyy9):

so the pythagorean theorem say that for a right triangle the square of hypothenuse c is equal the sum of square side a and square side b so than this mean in this case C^2 = A^2 +B^2 C^2 = 9 +81 C^2 = 90 C = sqrt 90 = sqrt(9*10) = 3sqrt10 C = 3sqrt10 hope this will help you

OpenStudy (anonymous):

thank you

jhonyy9 (jhonyy9):

was my pleasure